# LitE-SNN: Designing Lightweight and Efficient Spiking Neural Network through Spatial-Temporal Compressive Network Search and Joint Optimization

### v1

@@ Abstract
Spiking Neural Networks (SNNs) mimic the information-processing mechanisms of the human brain and are highly energy-efficient, making them well-suited for low-power edge devices. However, the pursuit of accuracy in current studies leads to large, long-timestep SNNs, conflicting with the resource constraints of these devices. In order to design lightweight and efficient SNNs, we propose a new approach named LitESNN that incorporates both spatial and temporal compression into the automated network design process. Spatially, we present a novel Compressive Convolution block (CompConv) to expand the search space to support pruning and mixed-precision quantization while utilizing the shared weights and pruning mask to reduce the computation. Temporally, we are the first to propose a compressive timestep search to identify the optimal number of timesteps under specific computation cost constraints. Finally, we formulate a joint optimization to simultaneously learn the architecture parameters and spatial-temporal compression strategies to achieve high performance while minimizing memory and computation costs. Experimental results on CIFAR10, CIFAR100, and Google Speech Command datasets demonstrate our proposed LitESNNs can achieve competitive or even higher accuracy with remarkably smaller model sizes and fewer computation costs. Furthermore, we validate the effectiveness of our LitESNN on the trade-off between accuracy and resource cost and show the superiority of our joint optimization. Additionally, we conduct energy analysis to further confirm the energy efficiency of LitESNN.
@@ 1 Introduction
Spiking Neural Networks (SNNs) have gained great attention in recent years due to their ability to mimic the information-processing mechanisms of the human brain. They use the timing of the signals (spikes) to communicate between neuronal units. A unit in SNNs is only active when it receives or emits a spike, enabling event-driven processing and high energy efficiency. Moreover, the received spikes are weight-accumulate into the membrane potential using only accumulate (AC) operations, which consumes much lower energy than the standard energy-intensive multiply-accumulate (MAC) strategy in artificial neural networks (ANNs) [?]. The processing mechanism of SNN can be realized in neuromorphic hardware, such as Truenorth [?], Loihi [?], etc.
@@
Neural architecture search (NAS) aims to automate the neural network design under specific resource constraints. Recent studies have applied NAS to SNN to enable flexible and effective network design and have demonstrated better performance compared to manually designed SNN architectures. However, most of these works do not consider resource constraints in the search process [??] and only aim at improving the accuracy. While [?] considers the number of spikes, thereby reducing communication overhead, this metric alone is insufficient for a comprehensive evaluation of resource consumption. To better evaluate resource utilization, we use model size for assessing memory footprint, which is also related to memory access energy, and computation complexity for measuring computational cost. These metrics offer more direct and comprehensive indicators of memory and computing resource demands [?].
@@
To design a lightweight and efficient SNN characterized by small model size and low computation complexity as well as high performance, we incorporate compression within the NAS framework. Existing model compression techniques, including pruning and quantization, reduce the number of parameters and the number of bits used to represent each parameter, respectively, benefiting memory footprint, power efficiency and computational complexity. However, existing SNN pruning and quantization works are all based on fixed architectures, remaining a question of how to integrate them into evolving SNN architectures in search process. In addition to spatial domain, SNNs uniquely feature a temporal domain that dictates time iterations (also known as timesteps). By compressing these iterations, we can directly reduce computational operations and save energy. As current SNN-based NAS methods all originate from CNNs and lack temporal considerations, the potential of implementing the compressive timesteps in NAS has not yet been exploited.
@@
In this paper, we introduce a novel approach named LitESNN that incorporates both spatial and temporal compression into the automated network design process, aimed at designing SNNs with compact models and lower computational complexity. Spatially, we propose a Compressive Convolution block (CompConv) that expands the search space to incorporate pruning and quantization. Considering the variable synaptic formats in neuromorphic hardware, CompConv supports mixed-precision quantization to make the network design more flexible and efficient. Meanwhile, we utilize the shared weights and shared pruning masks to reduce the computation introduced by mixed-precision. Temporally, we propose a compressive timestep search that is tailored for the temporal domain of SNNs. Given the computational cost constraint, the model can find the optimal timesteps in accordance with the architecture and spatial compression. Finally, we formulate a multi-objective joint optimization that simultaneously determines the network architecture and compression strategies. This joint optimization ensures a cohesive optimization of these factors, avoiding local optimal and suboptimal global solutions resulting from multiple single-objective optimizations. We conduct the experiments on two image datasets CIFAR10, CIFAR100 and a speech dataset Google Speech Command. Experimental results show that our proposed LitESNNs achieve competitive accuracy with remarkably smaller model sizes and fewer computation costs. Furthermore, we validate the effectiveness of our LitESNN in balancing accuracy with resource cost and the superiority of our joint optimization. Additionally, we conduct energy analysis to further confirm the energy efficiency of LitESNN.
@@ 2.1 Neural Architecture Search
Neural architecture search (NAS) aims to automatically design neural architectures that achieve optimal performance using limited resources. In artificial neural networks (ANNs), NAS networks have surpassed manually designed architectures on many tasks such as image classification [??]. In SNN domain, [?] presented a spike-aware NAS framework involving direct supernet training and an evolutionary search algorithm with spike-aware fitness. However, this approach is confined to exploring a few predefined blocks, limiting the potential to discover optimal designs beyond them. [?] selected the architecture that can represent diverse spike activation patterns across different data samples. [?] proposed a spike-based differentiable hierarchical search based on the DARTS [?] framework. The SNNs designed by these two works are more flexible and have better accuracy, however, their focus is solely on the accuracy aspect, neglecting critical factors like resource costs. Moreover, they do not consider the time domain during the search process, which is an important aspect of SNN.
@@ 2.2 Model Compression
Pruning involves removing unnecessary connections to reduce the number of parameters and computational needs. In ANNs, many existing studies employed important scores-based pruning [?], which involves training importance scores for each weight to quantify its relevance in predictions. In SNNs, [?] proposed a gradient-based rewiring method to learn the connectivity and weight parameters of SNNs. [?] addressed the pruning of deep SNNs by modeling the state transition of dendritic spine. [?] introduced the lottery ticket hypothesis-based pruning for sparse deep SNNs.
Quantization reduces the number of bits used to represent each parameter to reduce the memory footprint and computation complexity. Some ANN-based works [??] set the quantizer based on the weight and activation distribution. [?] converted the binary ANN to obtain a binary SNN. [?] used STDP to learn the one-bit weight synapses. However, these methods apply uniform bit-width, neglecting the different sensitivities of different filters. [?] employed a layer-wise Hessian trace analysis for bit-width allocation per layer, but it is limited to a three-layer shallow network.
There are several works that consider both pruning and quantization. [?] adapted ADMM optimization with spatio-temporal backpropagation for SNN pruning and quantization. However, this approach is performed sequentially and may lead to suboptimal results, e.g., the best network architecture for the dense and full-precision model is not necessarily the optimal one after pruning and quantization [?]. [?] presented a joint pruned and quantized SNN with STDP-based learning, but it is limited to a fixed and shallow (two layers) structure. Current SNN research lacks a comprehensive approach that simultaneously considers the neural architecture design, pruning and quantization. Therefore, we need a solution to jointly optimize these aspects. Similar work in ANN needs to train a large supernet and distill a number of smaller sub-networks, followed by training the accuracy predictors and evolutionary search [?]. The whole training process takes 100 GPU days on V100 GPU. Considering that SNNs usually require more training time than ANNs [?], it is not practical to utilize this approach for SNNs.
In time domain, [?] dynamically determines the number of timesteps during inference on an input-dependent basis. While this work effectively reduces timesteps, it is limited to the inference stage and cannot be optimized together with other training parameters, which also makes it unsuitable for use in NAS. In this paper, we will jointly optimize the pruning, mixed-precision quantization, and timestep compression through the NAS framework to develop lightweight and efficient SNNs.
@@ 3.1 Preliminary
We choose a widely used spiking neural model Leaky Integrate-and-Fire (LIF) to describe the neural behavior. The dynamics of iterative version [?] can be written as:
@@
where $\tau_{decay}$ denotes the membrane decay constant, $W_{ij}$ is the synaptic weight between the $j$-th neuron (or feature map) and the $i$-th neuron (or feature map). If the membrane potential at timestamp $t$ of the neuron in $n$-th layer $u^{t,n}$ is larger than the threshold $V_{th}$, the output spike $y_{i}^{t,n}$ is set to 1; otherwise it is set to 0. Our method builds upon the search space of spikeDHS [?], which is a spiking-based differentiable hierarchical search framework that can find best-performed architectures for SNNs. The search space consists of a number of cells and each cell (depicted in Figure 1(a)) is defined as a repeated and searchable unit with N nodes (depicted in Figure 1(b)), $\{x_{i}\}_{N}$. Each cell receives input from two previous cells and forms its output by concatenating all outputs of its nodes. Each node can be described by:
@@ 3.2 Compressive Convolution (CompConv)
To design lightweight and efficient SNNs, in this work, we incorporate compression techniques into the neural architecture search to take advantage of the efficiency and flexibility of automatic network design. To this end, we propose a novel compressive convolutional block (CompConv) to replace the standard convolution (green block in Figure 1) to enable the search of compressive SNNs. Neuromorphic hardware like Loihi is capable of supporting sparse kernels and variable synaptic formats [??], which inspired us to utilize weight pruning and mixed-precision quantization for compression in LitESNN.
@@
where $W^{(j)}$ represents the weight tensor of $j$-th branch, $Q_{j}$ is a quantization function that follows [?], which quantizes the input value by the $j$-th candidate bitwidth. $mask^{(j)}$ is a binary matrix where each element indicates whether the weights have been pruned (assigned a 0) or retained (assigned a 1) of $j$-th branch. The details of pruning method will be introduced in Supplementary. $W_{q}^{(j)}$ and $e^{(j)}(W)$ represent the quantized weights and operation of $j$-th branch respectively. $\beta^{(j)}$ is the weight of $j$-th operation and $B$ denotes the bit-width candidate space. However, this approach significantly increases the computation, as we need to update the weights and mask for each individual branch. Additionally, it has been found in [?] that branches with low $\beta$ values receive few gradients, potentially leading to the under-trained of weights and mask.
@@
Figure 2(c) illustrates the distributions of $W$, $\bar{e}(W)$, and $W_{output}$ with pruning rates. $W$ displays a symmetric, continuous distribution with a central peak diminishing towards the edges. $\bar{e}(W)$ represents the weighted average of $W$ quantized with 2-bit and 4-bit precision, showing a discrete but similarly trending distribution. The pruning rate curve indicates that weights with smaller absolute values undergo more pruning due to their lower importance. Thus, the distribution of the $W_{output}$ forms a pattern that is high on the sides and low at the center and far sides.
@@ 3.3 Compressive Timestep Search
To assess the computational operations associated with different timesteps, we follow the convention of the neuromorphic computing community to use the total synaptic operations (SynOps) metric [?], defined as:
@@
where $p$ is the pruning rate, $k_{h}$ and $k_{w}$ denote the kernel height and width, $C_{in}$ and $C_{out}$ denote the input and output channel size, $H$ and $W$ is the height and width of the feature map. $S$ is the average spike rate, which is originally defined as
@@
where $T$ is the total time length, $s^{(t)}$ indicates the average spike rate per neuron at the $t$-th timestep. We can see that the number of timesteps will affect the spike rate $S$ and, consequently, the computational cost of the network. As shown in Figure 2(d), to enable this search, we consider a range of candidate timestep count $\{1,2,\ldots,T\}$ and modify the Equation (13) by introducing trainable weights for each number of timesteps:
@@
where $\psi^{(t)}$ is the weight of having $t$ timesteps, $\frac{exp(\psi^{(t)})}{\sum_{t}^{T}exp(\psi^{(t)})}$ is the relative weight with the total sum equaling $1$. For illustration, consider a scenario where the total time length $T$ is 6, and the relative weights for each timestep $t\in\{1,2,3,4,5,6\}$ are trained as $[0,0,0,0,0.6,0.4]$. According to Equation (14), the spike rate $S$ would include the sum of spike rates for the first 5 timesteps, plus 40% of the spike rate from the $6$-th timestep. If the weight distribution changes to $[0,0,0,0,0,1]$, the spike rate $S$ would be computed as the sum of spike rates across all 6 timesteps, which degenerate to Equation (13). The trained weights $\psi$ allow us to evaluate the impact of the timestep selection on the overall network performance. We will introduce how to balance network performance against computational cost in Section 3.4.

### v2

@@ Abstract
Spiking Neural Networks (SNNs) mimic the information-processing mechanisms of the human brain and are highly energy-efficient, making them well-suited for low-power edge devices. However, the pursuit of accuracy in current studies leads to large, long-timestep SNNs, conflicting with the resource constraints of these devices. In order to design lightweight and efficient SNNs, we propose a new approach named LitE-SNN that incorporates both spatial and temporal compression into the automated network design process. Spatially, we present a novel Compressive Convolution block (CompConv) to expand the search space to support pruning and mixed-precision quantization. Temporally, we are the first to propose a compressive timestep search to identify the optimal number of timesteps under specific computation cost constraints. Finally, we formulate a joint optimization to simultaneously learn the architecture parameters and spatial-temporal compression strategies to achieve high performance while minimizing memory and computation costs. Experimental results on CIFAR-10, CIFAR-100, and Google Speech Command datasets demonstrate our proposed LitE-SNNs can achieve competitive or even higher accuracy with remarkably smaller model sizes and fewer computation costs.
@@ 1 Introduction
Spiking Neural Networks (SNNs) have gained great attention in recent years Liu et al. (2020); Zhang et al. (2021); Hu et al. (2021, 2023); Wang et al. (2023); Wei et al. (2024) due to their ability to mimic the information-processing mechanisms of the human brain. They use the timing of the signals (spikes) to communicate between neuronal units. A unit in SNNs is only active when it receives or emits a spike, enabling event-driven processing and high energy efficiency. The received spikes are weight-accumulated into the membrane potential using only accumulate (AC) operations, which consume much lower energy than the standard energy-intensive multiply-accumulate (MAC) strategy in Artificial Neural Networks (ANNs) Farabet et al. (2012). The processing mechanism of SNN can be realized in neuromorphic hardware, such as Truenorth Merolla et al. (2014), Loihi Davies et al. (2018), Darwin Ma et al. (2024), etc.
@@
Neural architecture search (NAS) aims to automate the neural network design under specific resource constraints. Recent studies have applied NAS to SNN to enable flexible and effective network design and have demonstrated better performance compared to manually designed SNN architectures. However, most of these works do not consider resource constraints in the search process Na et al. (2022); Che et al. (2022) and only aim at improving the accuracy. While Kim et al. consider the number of spikes, thereby reducing communication overhead Kim et al. (2022a), this metric alone is insufficient for a comprehensive evaluation of resource consumption. To better evaluate resource utilization, we use model size for assessing memory footprint, which is also related to memory access energy, and computation complexity for measuring computational cost. These metrics offer more direct and comprehensive indicators of memory and computing resource demands Zhang et al. (2016).
@@
To design a lightweight and efficient SNN characterized by small model size and low computation complexity as well as high performance, we incorporate compression within the NAS framework. Existing model compression techniques, including pruning and quantization, reduce the number of parameters and the number of bits used to represent each parameter, respectively, benefiting memory footprint, power efficiency and computational complexity. However, existing SNN pruning and quantization works are all based on fixed architectures, leaving open the question of how to integrate them into evolving SNN architectures in search process. In addition to spatial domain, SNNs uniquely feature a temporal domain that decides time iterations (also known as timesteps). By compressing these iterations, we can directly reduce computational operations and save energy. As current SNN-based NAS methods all originate from ANNs and lack temporal considerations, the potential of implementing the compressive timesteps in NAS has not yet been exploited.
@@
In this paper, we introduce a novel approach named LitE-SNN that incorporates both spatial and temporal compression into the automated network design process, aimed at designing SNNs with compact models and lower computational complexity. Spatially, we propose a Compressive Convolution block (CompConv) that expands the search space to incorporate pruning and quantization. Considering the variable synaptic formats in neuromorphic hardware, CompConv supports mixed-precision quantization to make the network design more flexible and efficient. Meanwhile, we utilize the shared weights and shared pruning masks to reduce the computation introduced by mixed-precision. Temporally, we propose a compressive timestep search that is tailored for the temporal domain of SNNs. Given the computational cost constraint, the model can find the optimal timesteps in accordance with the architecture and spatial compression. Finally, we formulate a multi-objective joint optimization that simultaneously determines the network architecture and compression strategies. This joint optimization ensures a cohesive optimization of these factors, avoiding local optimal and suboptimal global solutions resulting from multiple single-objective optimizations. We conduct the experiments on two image datasets CIFAR-10, CIFAR-100 and a speech dataset Google Speech Command. Experimental results show that our proposed LitE-SNNs achieve competitive accuracy with remarkably smaller model sizes and fewer computation costs. Furthermore, we validate the effectiveness of our LitE-SNNs in balancing accuracy with resource cost and the superiority of our joint optimization. Finally, we conduct energy analysis to further confirm the energy efficiency of LitE-SNNs.
@@ 2 Related Work
This section reviews current developments in neural architecture search (NAS) and model compression. NAS aims to automatically design neural architectures that achieve optimal performance using limited resources. Model compression aims to achieve lightweight neural networks with comparable performance, including pruning and quantization, etc.
@@ 2.1 Neural Architecture Search
In ANNs, NAS networks have surpassed manually designed architectures on many tasks such as image classification Liu et al. (2019); Cai et al. (2019). In the SNN domain, Na et al. (2022) presented a spike-aware NAS framework involving direct supernet training and an evolutionary search algorithm with spike-aware fitness. However, this approach is confined to exploring a few predefined blocks, limiting the potential to discover optimal designs beyond them. Kim et al. (2022a) selected the architecture that can represent diverse spike activation patterns across different data samples. Che et al. (2022) proposed a spike-based differentiable hierarchical search based on the DARTS Liu et al. (2019) framework. The SNNs designed by these two works are more flexible and have better accuracy, however, their focus is solely on the accuracy aspect, neglecting critical factors like resource costs. Moreover, they do not consider the time domain during the search process, which is an important aspect of SNN.
@@ 2.2 Model Compression
Pruning involves removing unnecessary connections to reduce the number of parameters and computational needs. In ANNs, many existing studies employed important scores-based pruning Sehwag et al. (2020), which involves training importance scores for each weight and masking less important weights. In SNNs, Chen et al. (2021) proposed a gradient-based rewiring method to learn the connectivity and weight parameters of SNNs. Chen et al. (2022) addressed the pruning of deep SNNs by modeling the state transition of dendritic spine. Kim et al. (2022b) introduced the lottery ticket hypothesis-based pruning for sparse deep SNNs.
Quantization reduces the number of bits used to represent each parameter to reduce the memory footprint and computation complexity. Some ANN-based works Cai et al. (2017); Cai and Vasconcelos (2020) set the quantizer based on the weight and activation distribution. Rueckauer et al. (2017) converted the binary ANN to obtain a binary SNN. Srinivasan and Roy (2019) used STDP to learn the one-bit weight synapses. However, these SNN methods apply uniform bit-width, neglecting the different sensitivities of different filters. Lui and Neftci (2021) employed a layer-wise Hessian trace analysis for mixed-precision quantization, allocating the bit-width per layer, but limited to a three-layer shallow network.
There are several works that consider both pruning and quantization. Deng et al. (2021) adapted alternating direction method of multipliers (ADMM) optimization with spatio-temporal backpropagation for SNN pruning and quantization. However, this approach is performed sequentially and may lead to suboptimal results, e.g., the best network architecture for the dense and full-precision model is not necessarily the optimal one after pruning and quantization Wang et al. (2020). Rathi et al. (2018) presented a joint pruned and quantized SNN with STDP-based learning, but it is limited to a fixed and shallow (two layers) structure. Current SNN research lacks a comprehensive approach that simultaneously considers the neural architecture design, pruning and quantization. Therefore, we need a solution to jointly optimize these aspects. Similar work in ANN needs to train a large supernet and distill a number of smaller sub-networks, followed by training the accuracy predictors and evolutionary search Wang et al. (2020). The whole training process takes 100 GPU days on V100 GPU. Considering that SNNs usually require more training time than ANNs Kim et al. (2022a), it is not practical to utilize this approach for SNNs.
In the time domain, Li et al. (2023) dynamically determines the number of timesteps during inference on an input-dependent basis. While this work effectively reduces timesteps, it is limited to the inference stage and cannot be optimized together with other training parameters, which also makes it unsuitable for use in NAS. In this paper, we will jointly optimize the pruning, mixed-precision quantization, and timestep compression through the NAS framework.
@@ 3 Method
In this section, we first introduce the preliminary knowledge of spiking neurons and architecture search space, then present our proposed spatial and temporal compression search, and finally explain the spatial-temporal joint optimization.
@@ 3.1 Preliminary
We choose a widely used spiking neural model Leaky Integrate-and-Fire (LIF) Zheng et al. (2021) to describe the neural behavior, which can be written as:
@@
where $\tau_{decay}$ denotes the membrane decay constant, $W_{ij}$ is the synaptic weight between the $j$-th neuron (or feature map) and the $i$-th neuron (or feature map). If the membrane potential at timestamp $t$ of the neuron in $n$-th layer $u^{t,n}$ is larger than the threshold $V_{th}$, the output spike $y_{i}^{t,n}$ is set to 1; otherwise it is set to 0. Our method builds upon the search space of spikeDHS Che et al. (2022), which is a spiking-based differentiable hierarchical search framework that can find best-performed architectures for SNNs. The search space consists of a number of cells and each cell (depicted in Figure 1(a)) is defined as a repeated and searchable unit with N nodes (depicted in Figure 1(b)), $\{x_{i}\}_{N}$. Each cell receives input from two previous cells and forms its output by concatenating all outputs of its nodes. Each node can be described by:
@@ 3.2 Compressive Convolution (CompConv)
To design lightweight and efficient SNNs, in this work, we incorporate compression techniques into the neural architecture search to take advantage of the efficiency and flexibility of automatic network design. To this end, we propose a novel compressive convolutional block (CompConv) to replace the standard convolution (green block in Figure 1) to enable the search of compressive SNNs. Neuromorphic hardware like Loihi is capable of supporting sparse kernels and variable synaptic formats Davies et al. (2018); Davies and others (2021), which inspired us to utilize weight pruning and mixed-precision quantization for compression.
@@
where $W^{(j)}$ represents the weight tensor of $j$-th branch, $Q_{j}$ is a quantization function that follows Cai et al. (2017), which quantizes the input value by the $j$-th candidate bitwidth. $mask^{(j)}$ is a binary matrix where each element indicates whether the weights have been pruned (assigned a 0) or retained (assigned a 1) of $j$-th branch. The pruning method follows the Sehwag et al. (2020). $W_{q}^{(j)}$ and $e^{(j)}(W)$ represent the quantized weights and operation of $j$-th branch respectively. $\beta^{(j)}$ is the weight of $j$-th operation and $B$ denotes the bit-width candidate space. However, this approach significantly increases the computation, as we need to update the weights and mask for each individual branch. Additionally, it has been found in Cai and Vasconcelos (2020) that branches with low $\beta$ values receive few gradients, potentially leading to the under-trained of weights and mask.
@@
Figure 2(c) illustrates the distributions of $W$, $\bar{e}(W)$, and $W_{output}$ with pruning. $W$ displays a symmetric, continuous distribution with a diminishing central peak. $\bar{e}(W)$, quantized at 2-bit and 4-bit precision, follows a similar discrete pattern. The pruning rate curve indicates that weights with smaller absolute values are pruned more for their lower importance, resulting in $W_{output}$ having higher values on the sides and lower at the center and far sides.
@@ 3.3 Compressive Timestep Search
To assess the computational operations associated with different timesteps, we follow the convention of the neuromorphic computing community to use the total synaptic operations (SynOps) metric Wu et al. (2021), defined as:
@@
where $p$ is the pruning rate, $k_{h}$ and $k_{w}$ are the kernel height and width, $C_{in}$ and $C_{out}$ the input and output channel sizes, $H$ and $W$ the height and width of the feature map. $S$ is the average spike rate, which is originally defined as
@@
where $T$ is the total time length, $s^{(t)}$ indicates the average spike rate per neuron at the $t$-th timestep. We observe that the number of timesteps affects the spike rate $S$ and, consequently, the network’s computational cost. As shown in Figure 2(d), to enable this search, we consider a range of candidate timestep counts $\{1,2,\ldots,T\}$ and modify the Equation (13) by introducing trainable weights for each timestep count:
@@
where $\psi^{(t)}$ is the weight of having $t$ timesteps, $\frac{exp(\psi^{(t)})}{\sum_{t}^{T}exp(\psi^{(t)})}$ is the relative weight with the total sum equaling $1$. For illustration, consider a scenario where the total time length $T$ is 6 with relative weights for each timestep $t\in\{1,2,3,4,5,6\}$ as $[0,0,0,0,0.6,0.4]$. Here, the spike rate $S$ includes the sum of spike rates for the first 5 timesteps plus 40% from the 6th. If weights change to $[0,0,0,0,0,1]$, $S$ equals the sum across all timesteps, matching Equation (13). The trained weights $\psi$ evaluate the impact of timestep selection on network performance. We will introduce how to balance network performance against computational cost in Section 3.4